I need to find the intersection points of a circle and an ellipse.
Circle: x^2 + (y-9)^2 = 16
Ellipse: x^2/16 + (y-1)^2/49 = 1
Please help me out. This is for a project, it is not in a book.

Answer by Alan3354(69443)

Circle: x^2 + (y-9)^2 = 16
Ellipse: x^2/16 + (y-1)^2/49 = 1
---------------
Ellipse x 16: x^2 + (16/49)*(y-1)^2 = 16
Set the 2 eqns equal to each other, solve for y.
--------------
y = 67/11 (Ignore y = 59/3)
---------------
Sub into original eqn
x = +- sqrt(912)/11 = ~ 2.74539
(-2.745,6.09),(+2.745,6.09)
-------------
email via the thank you note if you want additional help, or to see the graph.